# Fast, accurate and lightweight sequential simulation-based inference using Gaussian locally linear mappings

### 12 Mar 2024 v1

@@ 5.5 Results summary
In Table 2 we give a summary of the main results from all experiments where an objective comparison with NN-based methods was possible, including those in Supplementary Material. Instead, for the Lotka-Volterra model we did not run NN-based methods and we reported comparisons with SMC-ABC both in section 5.3 and in the Supplementary Material. Generally SeMPLE returns accurate inference, either by producing the smallest metric value (the smaller the better), as for Two Moons and Hyperboloid, or being second best after SNL, as for Bernoulli GLM and Ornstein-Uhlenbeck, also considering the larger variation of the SNPE-C results. However, even when SNL produces a smaller Wasserstein distance, it has to be balanced with respect to the running time. For Bernoulli GLM, SNL is between 18 and 75 times slower than SeMPLE, depending on the number of rounds, as SNL has to perform an expensive neural networks training at each round. For the same reason, when looking at Bernoulli GLM, SNPE-C is between 8 and 26 times slower than SeMPLE, again depending on how many times it needs to be retrained. When considering the Ornstein-Uhlenbeck model, the Wasserstein distances have a good degree of overlap between SNL and SeMPLE, but SNL is between 9.5 and 10.5 times slower than SeMPLE, while SNPE-C is between 17 and 22 times slower than SeMPLE. In Table 2 we highlight whether an experiment has a unimodal or a multimodal target, as this determines the type of metric we report. Namely we use the C2ST metric for multimodal targets and the Wasserstein distance for unimodal targets. The reasons for this are now discussed, however notice that in the Supplementary Material we report results for both metrics in each experiment. In our experiments, we found the C2ST metric (advocated in Lueckmann et al., 2021) useful to evaluate the inference quality with multimodal targets, when draws from a reference posterior are available. However, we also found that at times C2ST produced contradicting results with unimodal targets. An example is the Binomial GLM model considered in Supplementary Section 7. There, it is obvious that the marginal posteriors returned by SeMPLE are accurate and that inference improved between the first two rounds, however C2ST showed that inference was apparently deteriorating for SeMPLE from round $r=1$ to $r=2$, which was not the case in reality. We also computed Wasserstein distances, which could confirm that the inference quality via SeMPLE did improve from round 1 to round 2. However, for multimodal targets, it is true that Wasserstein distances may not be a suitable tool, see e.g. Balaji et al. (2019). We therefore warn the reader from blindly relying on a single performance metric. It is also important to examine marginal and joint posterior densities.
Table 2: Median metric results, across 10 independent runs (brackets include minimum and maximum values), and median runtime for posterior inference. Best (lowest) values are in bold. Notice, “Metric (final)” reports the metric at the final inference round, instead “#Model sims” gives the total number of model simulations across all rounds for a single run, and similarly for “Runtime”. SNPE-C and SNL used the default SBIBM setup except when the number of algorithm iterations is reduced from 10. The C2ST metric is reported for multimodal posterior distributions and the Wasserstein distance for unimodal posterior distributions according to the discussion in Section 5.5.
@@ 6 Conclusion and perspectives
Of course, SNL and SNPE-C could certainly be improved for each specific case study, but we did not examine or fine-tune SNL and SNPE-C on the many other possible configurations, as this can be influenced by many factors. While examining several case studies implemented in SBIBM, we restricted to simulation settings provided in the SBIBM package, including reusing the same NN architectures and training setups, see Supplementary Section 1 for details.

### 16 Mar 2024 v2

@@ 5.5 Results summary
Table 2: Median metric results, across 10 independent runs (brackets include minimum and maximum values), and median runtime for posterior inference. Best (lowest) values are in bold. Notice, “Metric (final)” reports the metric at the final inference round, instead “#Model sims” gives the total number of model simulations across all rounds for a single run, and similarly for “Runtime”. SNPE-C and SNL used the default SBIBM setup except when the number of algorithm rounds is reduced from 10. The C2ST metric is reported for multimodal posterior distributions and the Wasserstein distance for unimodal posterior distributions according to the discussion in Section 5.5.
@@
In Table 2 we give a summary of the main results from all experiments where an objective comparison with NN-based methods was possible, including those in the Appendices. Instead, for the Lotka-Volterra model we did not run NN-based methods and we reported comparisons with SMC-ABC both in section 5.3 and in Appendix I. Generally SeMPLE returns accurate inference, either by producing the smallest metric value (the smaller the better), as for Two Moons and Hyperboloid, or being second best after SNL, as for Bernoulli GLM and Ornstein-Uhlenbeck, also considering the larger variation of the SNPE-C results. However, even when SNL produces a smaller Wasserstein distance, it has to be balanced with respect to the running time. For Bernoulli GLM, SNL is between 18 and 75 times slower than SeMPLE, depending on the number of rounds, as SNL has to perform an expensive neural networks training at each round. For the same reason, when looking at Bernoulli GLM, SNPE-C is between 8 and 26 times slower than SeMPLE, again depending on how many times it needs to be retrained. When considering the Ornstein-Uhlenbeck model, the Wasserstein distances have a good degree of overlap between SNL and SeMPLE, but SNL is between 9.5 and 10.5 times slower than SeMPLE, while SNPE-C is between 17 and 22 times slower than SeMPLE. In Table 2 we highlight whether an experiment has a unimodal or a multimodal target, as this determines the type of metric we report. Namely we use the C2ST metric for multimodal targets and the Wasserstein distance for unimodal targets. The reasons for this are now discussed, however notice that in the Appendices we report results for both metrics in each experiment. In our experiments, we found the C2ST metric (advocated in Lueckmann et al., 2021) useful to evaluate the inference quality with multimodal targets, when draws from a reference posterior are available. However, we also found that at times C2ST produced contradicting results with unimodal targets. An example is the Binomial GLM model considered in Appendix G. There, it is obvious that the marginal posteriors returned by SeMPLE are accurate and that inference improved between the first two rounds, however C2ST showed that inference was apparently deteriorating for SeMPLE from round $r=1$ to $r=2$, which was not the case in reality. We also computed Wasserstein distances, which could confirm that the inference quality via SeMPLE did improve from round 1 to round 2. However, for multimodal targets, it is true that Wasserstein distances may not be a suitable tool, see e.g. Balaji et al. (2019). We therefore warn the reader from blindly relying on a single performance metric. It is also important to examine marginal and joint posterior densities.
@@ 6 Conclusion and perspectives
Of course, SNL and SNPE-C could certainly be improved for each specific case study, but we did not examine or fine-tune SNL and SNPE-C on the many other possible configurations, as this can be influenced by many factors. While examining several case studies implemented in SBIBM, we restricted to simulation settings provided in the SBIBM package, including reusing the same NN architectures and training setups, see Appendix A for details.